OpenStudy (kohai):

Verify: csc^2xsecx = secx + cscxcotx

OpenStudy (nincompoop):

can you type in the right hand side properly?

OpenStudy (kohai):

It is?

OpenStudy (nincompoop):

is 2xsecx all in the exponent form?

OpenStudy (kohai):

\[\csc^2xsecx = secx + cscxcotx\]

OpenStudy (nincompoop):

can you express in sine cosine?

OpenStudy (kohai):

\[(1+\cot^2)secx = \frac{ \cos^2x }{ \sin^2x } \times \frac{ 1 }{ cosx }\]

OpenStudy (nincompoop):

how about consistently including the left side?

OpenStudy (nincompoop):

cotangent as cosine/sine, cosecant as 1/sine and secant as 1/cosine

OpenStudy (kohai):

\[(1+\frac{ \cos^2x }{ \sin^2x })secx = secx + \frac{ \cos^2x }{ \sin^2x } \times \frac{ 1 }{ cosx }\]

OpenStudy (nincompoop):

\[\frac{ 1 }{ \cos (x) } \left( \frac{ 1 }{ \sin (x) } \right)^2= \frac{ 1 }{ \cos(x) }+\frac{ 1 }{ \sin (x) }\frac{ \cos (x) }{ \sin(x) }\]

OpenStudy (kohai):

How did you get that?

OpenStudy (nincompoop):

trig identity

OpenStudy (kohai):

I understand secx -> 1/cosx, but not the 1/sinx^2

OpenStudy (nincompoop):

csc = 1/sin -> csc^2 = (1/sin)^2

OpenStudy (kohai):

Oh, right. I see. Thanks!

OpenStudy (kohai):

And then the rest simplifies out

OpenStudy (nincompoop):

I made a booboo laughing out loud

OpenStudy (kohai):

\[\frac{ cosx }{ sinx^2 } = \frac{ 1 }{ cosxsinx^2 }\]

OpenStudy (nincompoop):

\[\frac{ 1}{ \cos x \sin x^2 }=\frac{ cosx^2 + \sin x^2}{ \cos x \sin x^2 }\]

OpenStudy (kohai):

\[\frac{ 1 }{ cosxsinx^2 } = \frac{ 1 }{ cosxsinx^2 }\]

OpenStudy (nincompoop):

now the denominator looks the same

OpenStudy (kohai):

And the numerator simplifies to 1 :)

OpenStudy (nincompoop):

you can cancel it out and then use pythagorean sinx^2+cos x^2 = 1 1 = 1

OpenStudy (kohai):

Awesome, thanks :D

OpenStudy (nincompoop):

np